a) Problem 1: What is your total for Inventory? See circle on the printed exercise.

b) Problem 1: What is your total for Net Income? See circle.

c) Problem 2: What is your total for Cost of Goods Sold? See circle.

d) Problem 2: What is your total for Net Income? See circle.

e) Problem 2: What is your total for Ending Inventory (in dollars)?

Instructions: Use the horizontal analysis on the next page to record the following transactions. Remember, any costs that you determine to bea product cost should be added to the inventory column. Period costs should be expensed. See the Self Study Review Problem at the end of the text chapter for guidance if necessary. Problem 1: Klyn Manufacturing Company experienced the following accounting events during its first year of operation. Except for the depreciation adjusting entries, all transactions are cash transactions. 1. Acquired $50,000 cash from the issue of common stock. 2. Paid $6,800 for the materials that were used to make its products. All products started were completed during the period. 3. Paid salaries of $4,300 to selling and administrative employees 4. Paid wages of $7,200 to production workers 5. On January 1, paid $9,000 to buy furniture used in selling and administrative offices. 6. Recorded annual depreciation on the furniture referred to in Event 5. The furniture had a $1,000 estimated salvage value and a 5-year useful life Use Straight-Line Depreciation. On January 1, paid $23,000 to buy manufacturing equipment. 7. 8. Recorded annual depreciation on the equipment referred to in Event 7. It had a $3,000 estimated salvage value and a 4-year useful life. Use Straight-Line Depreciation. Completed 4,000 units of product. Record the sale of 3,000 units of product. Determine the cost per unit and the sales price per unit assuming the sales price is "cost plus 60% of cost". 9. 10. Code the recognition of cost of goods sold for the units sold in Event 9
Problem 1 Continued: Show how these events would affect the balance sheet and income statement by recording them in a horizontal financial statements model like the one shown below. Make sure that the accounting equation balances after each transaction! Balance Sheet Income Statement Equi Assets Manf Equip Office Furn Common Net Inc Liab.+Stock Ea Cash Inventor Rev 1) 2) 3) 4) 5) 6) 7) 8) 9) 10) TOT
Problem 2: In its first year of operations, Baker Inc. paid $11,000 for direct materials used in its production facilities. Baker also paid $8,000 for factory workers' wages. Depreciation and rental payments related to the production factory amounted to $5,000. Baker also incurred general and administrative expenses in the amount of $11,000. The company produced 8,000 units and sold 6,000 units at a price of S9.00 a unit. What is the Cost of Goods Sold for the year? What is the Net Income for the year? What value would be assigned to Ending Inventory?
Answer #1
Computation of Selling price
Material 6800
Wages 7200
PRIME COST 14000

COGS PER UNIT:

14000/4000 = 3.5

Salary 4300
18300
Depreciation 6600

Equipment: (23-3)/4=5

Furniture: (9-1)/5=1.6

Total Cost of Production 24900
Cost per Unit 6.225 24900/4000
Selling price 9.96 6.225*160/100
So, Total Net Income 11205
Cogs 10500

2) COGS (11000+8000)/8000*6000=14250

NET INCOME (9*6000)-14250-16000=23750

CLOSING INVRNTORY: 4750